OpenStudy (anonymous):

A line contains the points (8, 9) and (–12, –7). Using point-slope form, write the equation of the line that is parallel to the given line and that passes through (–5, –15)

OpenStudy (anonymous):

first, calculate the slope of the line through (8, 9) and (–12, –7)... what is it?

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

good... so 4/5 is the slope of the line you want AND it needs to go through the point (-5,-15)... plug this into the point-slope form and what u got???

OpenStudy (anonymous):

HINT: point-slope form: \(\large y-y_0=m(x-x_0) \), where \(\large (x_0,y_0) \) is the given point.

OpenStudy (anonymous):

y+15=4/5(x+5)

OpenStudy (anonymous):

yep... good work.

OpenStudy (anonymous):

thank you for helping me out :)